Evaluate the following;

a. 9p2

b. 6p4

c. 12p3

d. 10p8


1
Expert's answer
2021-11-25T01:41:59-0500

"a. \frac{9!}{\left(9-2\right)!}=72\\

b. \frac{6!}{\left(6-4\right)!}=360\\

c. \frac{12!}{\left(12-3\right)!}=1320\\

d. \frac{10!}{\left(10-8\right)!}=1814400"